3x+2=11 what's the value of x
Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

3x+2=11
----------
Subtract 2
Divide by 3

Answer by jim_thompson5910(35256)

3x+2=11

3x+2-2=11-2

3x = 9

3x/3 = 9/3

x = 3


So the solution is x = 3